OpenStudy (anonymous):

The time it takes to train a newly hired telephone sales representative is normally distributed with a mean of 136 hours and a standard deviation of 8.9 hours. What is the probability that training for a given individual will last between 130 and 140 hours? Would a good starting point for this question be: P(130

OpenStudy (amistre64):

yes, now we could do a change of variable if need be

OpenStudy (amistre64):

are you able to use a ti83 or similar?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then all we need is a function called normalcdf

OpenStudy (amistre64):

normalcdf(low, high, mean, sd)

OpenStudy (amistre64):

2nd, vars, pick the normalcdf function and input the values

OpenStudy (anonymous):

how do you input the 140?

OpenStudy (anonymous):

And if you will have show work on test i will have to work z score...

OpenStudy (amistre64):

are you allowed to use a ti83 on a test?

OpenStudy (amistre64):

on my tests, we simply defined the function we used: normalcdf(130,140,136,8.9) = .4233

OpenStudy (anonymous):

test you can use any calculator, but the thing is the teacher expected work wrote out in steps...like calculating the z score, drawing the bell and writing the information in the bell...

OpenStudy (amistre64):

if we do a change of variable, then x = mean + z(sd)

OpenStudy (amistre64):

P(130<x<140) P(130 <mean + z(sd) < 140) P(130-mean < z(sd) < 140-mean) P((130-mean)/sd < z < (140-mean)/sd )

OpenStudy (anonymous):

I've never actually seen this you are showing me, so what the have us do is z=x-u/o...and i was thinking that you would find both z scores and subtract them to get the answer???

OpenStudy (amistre64):

subtracting z scores does nothing but tell us how many z scores fit between the interval. but we dont have a uniform distribution so 3 sd between a and b, is going to have a different value than 3 sd between c and d

OpenStudy (amistre64):

z scores allow us to look up tables that tell us the amount of area (probability) associted with it

OpenStudy (amistre64):

its those areas that we will work with

OpenStudy (amistre64):

but then tables are formatted by author preferences, so how we use a table will depend on your authors format

OpenStudy (anonymous):

Yes that is very true. But i see what you are showing me with the calculator but the study answers shows 0.4222 as the answer. Confused at moment

OpenStudy (amistre64):

.4233 is more exact, so it depends on where in the process you start rounding things and what you use to reference them with

OpenStudy (anonymous):

how did you enter the 130 and 140 values in calculator together?

OpenStudy (amistre64):

130-136 = 6/8.9 = 0.6741... = -0.67 140-136 = 4/8.9 = 0.4494... = 0.45 in the ti83 we have to get to the distribution menu: 2nd, VARS, then choose the normalcdf that prompts you with: normalcdf( input: 130, 140, 136, 8.9) and hit enter

OpenStudy (amistre64):

the normalcdf function takes 4 parameters: low, high, mean, sd

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=normalcdf%28-.67%2C.45%2C0%2C1%29 notice that when we rnd the z scores to 2 decies, we get the .4222 solution

OpenStudy (anonymous):

yes I got z scores of-.067 for 130 and 0.44 for 140

OpenStudy (amistre64):

the calculator wont rnd them, they will use them more precisely, therefore the 'difference' in solutions

OpenStudy (anonymous):

so i looked at the link you send and i plugged in the my numbers and I'm seeing what you are saying, but the z score of 130=-.67 which on table is 0.2514 and 140=.44 which on table is 0.6700. So what do you do with all that info to get 0.4222/0.4233?

OpenStudy (anonymous):

so you have add the two scores together and then subtract 1 from sum of z scores

OpenStudy (amistre64):

your z scores are giving you left tail areas 

OpenStudy (amistre64):

b is to big, a is to small, the middle part of it is just b-a

OpenStudy (amistre64):

and your z for 140 should rnd to .45

OpenStudy (anonymous):

Ok now I believe i have it. 130=-0.67=0.2514 and 140=0.45=0.6736. So 0.6736-0.2514=0.4222

OpenStudy (amistre64):

seems reasonable to me

OpenStudy (anonymous):

Thanks for your help and patience i feel a little slow after this;)

OpenStudy (amistre64):

yw, good luck :)